a theater group made appearances in two cities. The hotel charge before tax in the second city was $1000 lower than in the first. The tax in the first city was 10%, and the tax in the second city was 6%. The total hotel tax paid for the two cities was $
420. How much was the hotel charge in each city before tax?

Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52775)

.
a theater group made appearances in two cities.
The hotel charge before tax in the second city was $1000 lower than in the first.
The tax in the first city was 10%, and the tax in the second city was 6%.
The total hotel tax paid for the two cities was $420.
How much was the hotel charge in each city before tax?
~~~~~~~~~~~~~~~~~~~

Let x be the hotel charge in the first city before tax, in dollars.
Then that in the second city is (x-1000) dollars.


Write the total hotel tax paid for the two cities

    0.1x + 0.06(x-1000) = 420  dollars.


Simplify and find x

    0.1x + 0.06x - 60 = 420

        0.16x         = 420 + 60 = 480

            x                    = 480/0.16 = 3000.


ANSWER.  $3000 in the first city and $2000 in the second city.


CHECK.  0.1*3000 + 0.06*2000 = 300 + 120 = 420  dollars.   ! correct !

Solved.



Answer by josgarithmetic(39616)

CITY           BASE RATE         TAX RATE           TAX
   1            c+1000           10%               0.1(c+1000)
   2            c                 6%               0.06c
 TOTAL                                            420

Understand what equation to create and solve....